I am not sure how to solve the following word problem. I have more difficulty answering word problems than number problems. Words seem to overwhelm me more than numbers do. Any guidance that you can offer on this subject and the provided math problem would be greatly appreciated.
Problem: In a triangle the two larger angles differ by 10°. The small angle is 50° less than the largest angle. Find the measure of each angle.

Answer by mananth(16949)

if one angle is x of the larger two
then other is x-10
third angle is x-50
The sum of angles of triangle = 180 deg
x+x-10+x-50=180
3x-60=180
3x=240
x= 80
If x = 80
the second angle = 8-10=70
third angle = 80-50=30